A ROOM MEASURES 3M BY 4M BY 7M.
THE WALLS NEED 2 COATS OF PAINT. A 4-L CAN OF PAINT COVERS 40M SQUARED.
HOW MUCH PAINT SHOULD TANYA BUY?
I KNOW THE ANSWER BUT NOT HOW I GOT IT.

Answer by benni1013(206)

When doing surface area find all the possible combination of the numbers and multiply it by two.
(3*4)+(3*7)+(7*4)=61*2=122 meters^2
This is your surface area for the room.
122/40=3 and a little so you need 4 but you need two coats so you need 7. The decimal when you originally divide is so negligible that you won't need 8 cans.
This tells you how many cans of paint she needs to buy.
7*4=28 liters of paint. 7 equals the number of cans to buy 4 equals the amount in each can.